# Text2midi: Generating Symbolic Music from Captions

### v1

@@ Introduction
  • Tempo Bin (TB): the ratio of files of which the extracted tempo (in terms of beats per minute (bpm)), falls within into the tempo bin extracted from the ground truth MIDI file, with bin borders defined at 40, 60, 70, 90, 110, 140, 160, 210 bpm. The tempo is extracted with music21 (Cuthbert and Ariza 2010).
@@
Table 2 shows the results of the objective evaluation. Evaluation metrics are computed on MIDI files generated (from the same caption) by both text2midi, MuseCoco (Lu et al. 2023), and the original MIDI files (ground truth) from the data set (MidiCaps).
text2midi MidiCaps MuseCoco
CR $\uparrow$ 2.14 3.43 2.12
CLAP $\uparrow$ 0.22 0.26 0.21
TB (%) $\uparrow$ 27.85 - 21.71
TBT (%) $\uparrow$ 57.78 - 54.63
CK (%) $\uparrow$ 7.69 - 13.70
CKD (%) $\uparrow$ 14.80 - 14.59
@@
We observe an average compression ratio for the generated examples by text2midi of 2.14 for the generated MIDI. Looking at the results by MuseCoco we notice the ratio is lower (2.12). This indicates text2midi even fares better in creating long-term structure and repeating pattern compared to MuseCoco.
@@
While the CLAP score provides a general indication of how well the text and MIDI align, we further analyze specific aspects in which the generated MIDI resembled the ground truth MIDI. It is important to note that as these metrics compare the generated MIDI with the ground truth MIDI, they cannot be computed for MidiCaps, which serves as the ground truth here. We show that in terms of both tempo and key attributes, text2midi outperforms the MuseCoco model. The tempo bin (TB) metric reaches 27.85 compared to 21.71 for MuseCoco, and the TBT reaches 57.78 compared to 54.63 for MuseCoco. These ratios indicate that about 58% of the tempo instructions are very close to the desired tempo. The CKD ratios for text2midi are also higher than MuseCoco (14.80 % vs 14.59%), indicating better key matching. We report results on all of MidiCaps test set on our github page  [^6] https://github.com/AMAAI-Lab/Text2midi. In terms of inference speed, we notice that text2midi is much faster compared to MuseCoco as a result of unifying the text-to-attribute and attribute-to-music models into a single holistic approach. Based on our observation, text2midi takes around 55 seconds to generate a 40-second long MIDI file, compared to 120 seconds for MuseCoco (inference performed on GPU in both cases). This comprehensive evaluation demonstrates that text2midi consistently surpasses MuseCoco on all objective metrics, despite training on a data set of only one fifth of the size used to train MuseCoco.

### v2

@@ Introduction
  • Tempo Bin (TB): the ratio of files of which the extracted tempo (in terms of beats per minute (bpm)), falls within the tempo bin extracted from the ground truth MIDI file, with bin borders defined at 40, 60, 70, 90, 110, 140, 160, 210 bpm. The tempo is extracted with music21 (Cuthbert and Ariza 2010).
@@
ble 2 shows the results of the objective evaluation. Evaluation metrics are computed on MIDI files generated (from the same caption) by both text2midi, MuseCoco (Lu et al. 2023), and the original MIDI files (ground truth) from the data set (MidiCaps).
Model MidiCaps MuseCoco P-Val
CR $\uparrow$ 2.31 3.43 2.12 $<0.0001$
CLAP $\uparrow$ 0.22 0.26 0.21 0.0102
TB (%) $\uparrow$ 39.70 - 21.71 0.1102
TBT (%) $\uparrow$ 65.80 - 54.63 0.2051
CK (%) $\uparrow$ 33.60 - 13.70 $<0.0001$
CKD (%) $\uparrow$ 35.60 - 14.59 $<0.0001$
@@
Generating music with long-term structure is a known challenge (Bhandari and Colton 2024). We use the compression ratio, as done by Chuan and Herremans (2018), to evaluate the presence of repeated patterns in the MIDI files. text2midi achieves an average compression ratio of 2.31, significantly higher than MuseCoco’s 2.12 ($p<0.0001$). This result demonstrates that text2midi outperforms MuseCoco in creating long-term structure and repeating patterns.
@@
While the CLAP score provides a general indication of how well the text and MIDI align, we further analyze specific aspects in which the generated MIDI resembled the ground truth MIDI. It is important to note that as these metrics compare the generated MIDI with the ground truth MIDI, they cannot be computed for MidiCaps, which serves as the ground truth here. We show that in terms of both tempo and key attributes, text2midi outperforms the MuseCoco model. The tempo bin (TB) metric reaches 39.70 compared to 21.71 for MuseCoco, and the TBT reaches 65.80 compared to 54.63 for MuseCoco. These ratios indicate that about 66% of the tempo instructions are very close to the desired tempo. The CKD ratios for text2midi are also higher than MuseCoco (35.60% vs 14.59%), indicating better key matching. We note that both CK and CKD are statistically significant ($p<0.0001$). In terms of inference speed, we notice that text2midi is much faster compared to MuseCoco as a result of unifying the text-to-attribute and attribute-to-music models into a single holistic approach. Based on our observation, text2midi takes around 55 seconds to generate a 40-second long MIDI file, compared to 120 seconds for MuseCoco (inference performed on GPU in both cases). This comprehensive evaluation demonstrates that text2midi consistently surpasses MuseCoco on all objective metrics, despite training on a data set of only one fifth of the size used to train MuseCoco.